# Strain stiffening due to stretching of entangled particles in random packings of granular materials

### v1

@@ Abstract
Stress-strain relations for random packings of entangling chains under triaxial compression can exhibit strain stiffening and sustain stresses several orders-of-magnitude beyond typical granular materials. X-ray tomography reveals the transition to this strong strain stiffening occurs when chains are long enough to entangle an average of about one chain each, which results in system-filling clusters of entangled chains, similar to the Erdös-Rényi model for randomly connected graphs. The number of entanglements is nearly proportional to the area surrounded by entangling particles with an excluded volume effect, thus the existence of system-filling clusters of entanglements can be predicted assuming random particle positions and orientations with an excluded volume effect if the particle shapes in the packing are known. A tendency was found for chain links to stretch when the packing was strained. The slope of the stress-strain relation of the packing can be calculated from a mean-field model consisting of the product of the effective extensional modulus of the chain, packing fraction, probability of stretched links, and the ratio of strain of stretched links to packing strain. The stress-strain model requires as input measurements of the ratio between local particle deformation and global average strain, and the probability of stretching for non-rigid particles. This results in a quadratic prediction for the stress-strain curve, with a curvature that agrees with experiments within the model uncertainties ($\leq 37\%$). This model explains that the strength of these packings comes from stretching of the links of chains, but only when the system-filling network of entanglements provides constraints that prevents failure by shear banding, so that particles must be deformed to move further under strain. In this model, the increasing slope of the stress-strain curve is mainly due to the fraction of stretched links increasing with strain. This model for the stress-strain relation is shown to be generalizable to different shapes of entangling particles by applying it to staples. The permanent plastic deformation of staples allows measuring statistical quantities like the probability of bending and average bend angle from inspection of a poured out sample after a triaxial compression, without the need for in-situ imaging.
@@ I Introduction
Studying the mechanics of randomly packed entangling granular materials may lead to understanding of the mechanics of a broader class of materials that can entangle. This includes random packings of rigid particles with concave regions which can interlock Karayiannis et al. (2009); Mewis and Wagner (2012). Such particles have been of interest, for example as a pourable construction material with high strength but without need for a binding material Danel (1953); Dierichs and Menges (2016). There are many strain stiffening systems consisting of entangled systems of flexible fibers, including collections of disorganized matted fibers like cotton balls Poquillon et al. (2005), some knitted textiles Weeger et al. (2018), and birds nests Weiner et al. (2020), where the mechanics responsible for strain stiffening are not yet understood. There are opportunities for unique control of mechanical properties in entangled systems. For example, entangled structures of hard particles can have bending stiffness that is tunable with applied pressure over a much higher range than traditional soils Wang et al. (2021). The ability to manufacture particles with arms that can be activated to control entanglements opens up more possibilities for controllable mechanics Savoie et al. (2023).
@@
Strain stiffening is known to occur in many polymeric materials as well, where entanglements can be spontaneously driven by thermal motion Treolar (1949). Models for strain stiffening in polymers rely on the thermodynamics of wormlike chains, and have been applied only to relatively weak strain stiffening in polymer systems Buckley and Jones (1995); Hoy and Robbins (2006); Vorselaars et al. (2009); Hoy and O’Hern (2010). The wormlike chain model has been adapted to athermal granular systems with localized strain using the mechanism of self-amplified friction Dumont et al. (2018), but we find it does not characterize the type of strain stiffening we have observed in triaxial compression tests of entangling granular materials (see Appendix A).
@@
It has been shown in 2-dimensional simulations of sheared granular piles that even a single chain can increase strength and produce strain stiffening by suppressing the usual failure mode of shear banding when the chain is placed so that it crosses the path of the shear band Regev and Reichhardt (2013). Suppression of shear banding is likely necessary for the strength to greatly exceed the external confining stress. This may be similar to the way that long fibers interspersed in random packings can suppress the formation of shear bands and increase strength in composite materials such as concrete, but typically the strength increase is not more than a factor of 2 Ashour et al. (1992). A stronger constraint seems to be required to explain the observation of stresses several orders of magnitude times the confining stress Brown et al. (2012).
@@
In the remainder of this paper we develop models to predict which particle shapes will lead to system-spanning clusters of entanglements and strong strain stiffening in random packings, and the corresponding stress-strain curve. Section II explains the methods for stress-strain measurements. Section III.1 reviews stress-strain curves from triaxial compression measurements on random packings of granular chains from our previous letter Brown et al. (2012), which will be used to motivate and test quantitative models. Section III.2 shows measurements of the area enclosed by chains from x-ray tomography measurements of random packings of granular chains. We use this as input to a model for the probability of another chain passing through this area and thus being entangled in a random packing, to determine whether there will be system-filling clusters of entanglements and thus strong strain stiffening in Sec. III.3. Section III.4 shows measurements of link stretching when the packing is under strain to determine whether strain stiffening relates to tension on links or compression of loops. In section III.5, we work out a quantitative prediction for the stress-strain curve based on the increased probability of chains being stretched at higher strain, and the amount of stretching. The result is compared to stress-strain measurements for chain packings. In Sec. IV, we show how this model for the stress-strain relation can be applied to another type of entangled granular material: rigid U-shaped staples. In the appendices, we include additional supporting calculations, structural measurements obtained from x-ray tomography measurements, and particle stiffness measurements.
@@ II Methods
We prepared packings of chains for triaxial compression experiments Lambe and Whitman (1969). For these measurements we used brass beads with diameter $a=2.06\pm 0.02$ mm and minimum loop circumference $\xi=8$ beads. Measurements of the stiffness of individual chains in response to tension and bending are shown in Appendices B and C. Each chain in a packing was cut to the same length $N$ in units of the number of beads. The chains were poured a few at a time under gravity into a flexible elastic membrane in a cylindrical shape with solid end caps on the top and bottom, as shown in Fig. 1b. The granular material was compacted by dropping the sample in the membrane attached to the bottom end cap from a height of 30 mm, for 10 repetitions. The flexible membrane stretches just enough to provide a confining stress to hold up the weight of the sample and endcap, and allows for radial expansion when the sample is strained. We report data for end caps with diameter 50.8 mm and initial sample heights approximately 50 mm. The heights vary by a few mm from run to run due to different levels of compaction (packing fraction values are shown in Appendix D). The chains can be seen in Fig. 1b to form tight loops within the packing, as small as the minimum loop circumference $\xi$ Zhou et al. (2009).
@@ III.1 Strong strain stiffening
In this subsection, we summarize results from our previous letter Brown et al. (2012). We show in Fig. 2 the stress-strain relation $\sigma(\gamma)$ for a packing of chains with length $N\approx 10^{4}$ beads, much longer than their minimum loop circumference $\xi=8$ beads. Strong strain stiffening can be seen as the region of positive curvature of $\sigma(\gamma)$. A striking feature is that the stress reaches the order of 10 MPa before the packing fails, which is about $10^{3}$ times the stress from the weight of the packing on the elastic membrane at the boundary which usually determines the maximum strength of packings of convex particles. The latter behavior is shown for $N=1$ (unlinked beads) in the inset of Fig. 2, where the kink at around $\gamma=0.05$ is on the order of the stress on the membrane from the weight of the packing and endcap, and the positive slope for $\gamma\stackrel{{\scriptstyle>}}{{{}_{\sim}}}0.05$ matches the contribution of the membrane stiffness as it is deformed. The maximum strength of the long chain packing reaches up to the order of 1/10 the strength per density of the solid brass that makes up the chains. During the compression of the long chains in Fig. 2 we started to hear chains break at the point where the maximum stress was reached, and after each measurement that exceeded this ultimate stress, we counted approximately 10% of the links in the chains to be broken. This breaking, and the ability to support stresses much greater than the confining stress suggests that the strength of the links are the limiting factor in the packing strength.
@@
To illustrate the transition from convex particle behavior in the limit of unlinked beads ($N=1$) to strain stiffening in the limit of large $N$, we show stress-strain curves for packings of chains with different lengths $N$ in Fig. 3. The data fall into two distinct bands: the upper band of data with chain lengths $N\geq 9$ exhibits strong strain stiffening, while the lower band of data with $N\leq 8$ is much weaker. Packings with chain lengths $5\leq N\leq 8$ exhibit very mild strain stiffening, while those with $N\leq 3$ exhibit the strain-softening typical of convex particles. While the stress $\sigma$ tends to get larger with increasing $N$, the curves in Fig. 3 are not monotonically increasing with $N$, indicating a large run-to-run variation ($\sim 30\%$) compared to any trends in $N$ for $N>8$.
@@ III.2 Entangled area
To develop a quantitative model to predict the probability of entanglements, we first measure the free enclosed area $A$ enclosed by each chain which is available for other chains to pass through. This is calculated from the x-ray tomography data as the area in the entanglement manifold that is at least 1 bead diameter away from a bead in the entangling chain, as illustrated in Fig. 4b. This condition accounts for the fact that the finite bead size excludes the center of other beads from overlapping with the beads in the entangling chain, a.k.a an excluded volume effect. The average values for each contour length $m$ are shown in Fig. 5, shown in units of bead diameter squared. The shorter contour lengths have very little free enclosed area $A$ because the chains are short compared to the persistence length or minimum loop size and tend to be straighter. For contour lengths of $m\leq 3$, the finite bead size completely excludes chains from passing through the enclosed area for this minimum loop size. This explains why no entanglements were observed for $m\leq 3$ Brown et al. (2012). Since chains that are longer can be bent into larger arcs, the free enclosed area tends to increase with $m$, similar to the increase in number of entanglements with $m$ Brown et al. (2012).
Figure 6: Number of entanglements $n$ per enclosure as a function of free enclosed area $A$. The points correspond to different contour lengths $m$, using the same data as Fig. 5. Circles: chain length $N=20$. Diamonds: $N=10$. Open symbols: $\gamma=0$. Filled symbols: $\gamma=0.2$. Crossed symbols: plotted against free enclosed area $A$. Non-crossed symbols: plotted against effective free enclosed area $A_{eff}$ including constraints in the third dimension. Lines: linear fits to $N=10$ data. The near proportionality between $n$ and $A$ or $A_{eff}$ suggests that the probability of entanglement is determined by random arrangements with excluded volume effects, so the average number of entanglements and whether or not there are likely to be system-filling clusters of entanglements can be predicted from particle shapes.
@@
The number of entanglements $n$ is predicted based on random arrangements with excluded volume effects using a model calculation that was originally developed for staple packings Gravish et al. (2012). It involves equating two calculations of volume of entangled particles inside the volume made by an entanglement manifold of thickness $\delta$ – in this case we set $\delta$ equal to one bead diameter $a$. The average volume of chains in an entangled volume $V_{ent}$ can be estimated as the average number of entanglements $n$ times the volume per unit length of chain taken up by an entangled bead $\pi a^{3}/6\bar{d}$ (with average center-to-center spacing $\bar{d}$) times the length $\delta/\cos\Theta$ in the entangled plane of thickness $\delta$, where $\Theta$ is the angle of the entangled chain relative to a line perpendicular to the manifold (see illustration in Fig. 22 of Appendix E). This results in $V_{ent}=n\pi a^{3}\delta/6\bar{d}\cos\Theta$. This is balanced with a prediction of $V_{ent}$ based on the assumption that particle positions are random so the volume of entangled particles is equal to the packing fraction times the free volume available inside the entanglement. This is calculated as $V_{ent}=\phi A(\Theta,\Phi)\delta$, where $A(\Theta,\Phi)$ is a modification of $A$ which also accounts for the three-dimensional excluded volume from different possible angles $\Theta$ and $\Phi$ around this axis. Equating the two calculations of $V_{ent}$ results in $n=6\bar{d}\phi\langle A(\Theta,\Phi)\cos\Theta\rangle/\pi a^{3}=0.93\langle A(\Theta,\Phi)\cos\Theta\rangle/a^{2}$ for $\phi=0.44\pm 0.03$ (see Appendix D for measurements of $\phi$) and average center-to-center distance $\bar{d}/a=1.11$ (measured from x-ray tomography). The effective free enclosed area $A_{eff}=\langle A(\Theta,\Phi)\cos\Theta\rangle$ accounts for excluded volume effects in three dimensions, as well as the larger cross sections of entangled chains at various angles $\Theta$. $A_{eff}/A$ corresponds to the parameter $1/\alpha$ introduced as a fit parameter by Gravish et al. (2012). The calculation of $A_{eff}$ from our x-ray tomography data is explained in Appendix E. The number of entanglements $n$ as a function of $A_{eff}$ calculated from x-ray tomography data is shown in Fig. 6. This also appears nearly linear, as there is nearly a constant ratio $A_{eff}/A=0.27\pm 0.01$ in the strong strain-stiffening range with $m\geq 8$ (see Appendix E). A linear fit yields $n=(0.87\pm 0.14)A_{eff}/a^{2}$ for $N=10$ with percentage errors, consistent with the predicted value of $n=0.93A_{eff}/a^{2}$. This confirms that the assumption of random positions and orientations with excluded volume effects can relate the number of entanglements $n$ to the free enclosed area $A_{eff}$ if the shape of the particles in the packing is known, at least for $N=10$ near the onset of strong strain stiffening. The measured ratio $n/A_{eff}$ decreases for $N=20$ and for increasing $m$, suggesting that longer chains have additional constraints that reduce the number of entanglements, for example due to the maximum bend angle of chains which puts constraints on particle positions in the same chain. Since $n/A$ is more linear than $n/A_{eff}$ for small $m$, the larger $n/A_{eff}$ for small $m$ with the assumption of random particle orientations $\Theta$ in $A_{eff}$ indicates the angles $\Theta$ are more perpendicular to the entangled plane than random orientations, which could be a result of contact forces tending to push entangled chains into the small range of angles that they will fit.

### v2

@@ Abstract
Stress-strain relations for random packings of entangling chains under triaxial compression can exhibit strain stiffening and sustain stresses several orders-of-magnitude beyond typical granular materials. X-ray tomography reveals the transition to this strong strain stiffening occurs when chains are long enough to entangle an average of about one chain each, which results in system-filling clusters of entangled chains, similar to the Erdös-Rényi model for randomly connected graphs. The number of entanglements is nearly proportional to the area surrounded by entangling particles with an excluded volume effect, thus the existence of system-filling clusters of entanglements can be predicted assuming random particle positions and orientations with an excluded volume effect if the particle shapes in the packing are known. A tendency was found for chain links to stretch when the packing was strained This suggests that the strength of these packings comes from stretching of the links of chains, but only when the system-filling network of entanglements provides constraints that prevents failure by shear banding, so that particles must be deformed to move further under strain. The slope of the stress-strain relation of a packing can be calculated from a mean-field model consisting of the product of the effective extensional modulus of the chain, packing fraction, probability of stretched links, and the ratio of strain of stretched links to packing strain. In this model, the increasing slope of the stress-strain curve is mainly due to the fraction of stretched links increasing with strain, and assuming the fraction of stretched links is proportional to strain results in a quadratic prediction for the stress-strain curve. The stress-strain model requires as input measurements of the ratio between local particle deformation and global average strain, and the probability of stretching for non-rigid particles, resulting in a quadratic curvature that agrees with experiments within the run-to-run variation ($30\%$). This model for the stress-strain relation is shown to be generalizable to different shapes of entangling particles by applying it to staples, where the packing strength comes from the bending of staples instead of stretching links. The permanent plastic deformation of staples allows measuring statistical quantities from inspection of a poured out sample after a triaxial compression, without the need for in-situ imaging. Both the probability of staples bending and the average bend angle of the arms were found to increase with strain, and these inputs into the model result in a quadratic curvature of the stress-strain that agrees with experiments within the model uncertainties ($37\%$).
@@ I Introduction
Studying the mechanics of randomly packed entangling granular materials may lead to understanding of the mechanics of a broader class of materials that can entangle. This includes random packings of rigid particles with concave regions which can interlock Karayiannis et al. (2009); Mewis and Wagner (2012). Such particles have been of interest, for example as a pourable construction material with high strength but without need for a binding material Danel (1953); Dierichs and Menges (2016). Strain stiffening is observed in athermal gels formed by semiflexible fibers, where it is interpreted as a critical scaling behavior beyond a critical strain where fiber stretching becomes the dominant source of stress van Oosten et al. (2016). There are many strain stiffening systems consisting of entangled systems of flexible fibers, including collections of disorganized matted fibers like cotton balls Poquillon et al. (2005), some knitted textiles Weeger et al. (2018), and birds nests Weiner et al. (2020), where the mechanics responsible for strain stiffening are not yet understood. There are opportunities for unique control of mechanical properties in entangled systems. For example, entangled structures of hard particles can have bending stiffness that is tunable with applied pressure over a much higher range than traditional soils Wang et al. (2021). The ability to manufacture particles with arms that can be activated to control entanglements opens up more possibilities for controllable mechanics Savoie et al. (2023).
@@
Strain stiffening is known to occur in many thermal polymeric materials as well, where entanglements can be spontaneously driven by thermal motion Treolar (1949). Models for strain stiffening in polymers usually rely on the thermodynamics of wormlike chains, and have been applied only to relatively weak strain stiffening in polymer systems Buckley and Jones (1995); Hoy and Robbins (2006); Vorselaars et al. (2009); Hoy and O’Hern (2010). The wormlike chain model has been adapted to athermal granular systems with localized strain using the mechanism of self-amplified friction Dumont et al. (2018), but we find it does not characterize the type of strain stiffening we have observed in triaxial compression tests of entangling granular materials (see Appendix A).
@@
It has been shown in 2-dimensional simulations of sheared granular piles that even a single chain can increase strength and produce strain stiffening by suppressing the usual failure mode of shear banding when the chain is placed so that it crosses the path of the shear band Regev and Reichhardt (2013). Suppression of shear banding is likely necessary for the strength to greatly exceed the external confining stress. This may be similar to the way that long fibers interspersed in random packings can suppress the formation of shear bands and increase strength in composite materials such as concrete, but typically the strength increase is not more than a factor of 2 AHFW92. A stronger constraint seems to be required to explain the observation of stresses several orders of magnitude times the confining stress Brown et al. (2012).
@@
In the remainder of this paper we develop models to predict which particle shapes will lead to system-spanning clusters of entanglements and strong strain stiffening in random packings, and the corresponding stress-strain curve. Section II explains the methods for stress-strain measurements. Section III.1 reviews stress-strain curves from triaxial compression measurements on random packings of granular chains from our previous letter Brown et al. (2012), which will be used to motivate and test quantitative models. Section III.2 shows measurements of the area enclosed by chains from x-ray tomography of random packings of granular chains. We use this as input to a model for the probability of another chain passing through this area and thus being entangled in a random packing, to determine whether there will be system-filling clusters of entanglements and thus strong strain stiffening in Sec. III.3. Section III.4 shows measurements of link stretching when the packing is under strain to determine whether strain stiffening relates to tension on links or compression of loops. In section III.5, we work out a quantitative prediction for the stress-strain curve based on the increased probability of chains being stretched at higher strain, and the amount of stretching. The result is compared to stress-strain measurements for chain packings. In Sec. IV, we show how this model for the stress-strain relation can be applied to another type of entangled granular material: rigid U-shaped staples. In the appendices, we include additional supporting calculations, structural measurements obtained from x-ray tomography measurements, and particle stiffness measurements.
@@ II Methods
We prepared packings of chains for triaxial compression experiments Lambe and Whitman (1969). For these measurements we used brass beads with diameter $a=2.06\pm 0.02$ mm (all uncertainties are presented as one standard error, corresponding to a 68% confidence interval) and minimum loop circumference $\xi=8$ beads. Measurements of the stiffness of individual chains in response to tension and bending are shown in Appendices B and C. Each chain in a packing was cut to the same length $N$ in units of the number of beads. The chains were poured a few at a time under gravity into a flexible elastic membrane in a cylindrical shape with solid end caps on the top and bottom, as shown in Fig. 1b. The granular material was compacted by dropping the sample in the membrane attached to the bottom end cap from a height of 30 mm, for 10 repetitions. The flexible membrane stretches just enough to provide a confining stress to hold up the weight of the sample and endcap, and allows for radial expansion when the sample is strained. We report data for end caps with diameter 50.8 mm and initial sample heights approximately 50 mm. The heights vary by a few mm from run to run due to different levels of compaction (packing fraction values are shown in Appendix D). The chains can be seen in Fig. 1b to form tight loops within the packing, as small as the minimum loop circumference $\xi$ Zhou et al. (2009).
@@ III.1 Strong strain stiffening
In this subsection, we summarize results from our previous letter Brown et al. (2012). We show in Fig. 2 the stress-strain relation $\sigma(\gamma)$ for a packing of chains with length $N\approx 10^{4}$ beads, much longer than their minimum loop circumference $\xi=8$ beads. Strong strain stiffening can be seen as the region of positive curvature of $\sigma(\gamma)$. A striking feature is that the stress reaches the order of 10 MPa before the packing fails, which is about $10^{3}$ times the stress from the weight of the packing pushing on the elastic membrane at the boundary which usually determines the scale of maximum strength of packings of convex particles. The latter behavior is shown for $N=1$ (unlinked beads) in the inset of Fig. 2, where the positive slope for $\gamma\stackrel{{\scriptstyle>}}{{{}_{\sim}}}0.05$ matches the contribution of the membrane stiffness as it is deformed by the radially expanding packing, which is steeper while the packing is initially dilating. In other words, the strength of unlinked bead packings is largely determined by the confining stress rather than representing an intrinsic strength of the packing. The maximum strength of the long chain packing reaches up to the order of 1/10 the strength per density of the solid brass that makes up the chains. During the compression of the long chains in Fig. 2 we started to hear chains break at the point where the maximum stress was reached, and after each measurement that exceeded this ultimate stress, we counted approximately 10% of the links in the chains to be broken. This breaking, and the ability to support stresses much greater than the confining stress suggests that the strength of the links are the limiting factor in the packing strength.
@@
To illustrate the transition from convex particle behavior in the limit of unlinked beads ($N=1$) to strain stiffening in the limit of large $N$, we show stress-strain curves for packings of chains with different lengths $N$ in Fig. 3. The data fall into two distinct bands: the upper band of data with chain lengths $N\geq 9$ exhibits strong strain stiffening, while the lower band of data with $N\leq 8$ is much weaker. Packings with chain lengths $5\leq N\leq 8$ exhibit mild strain stiffening, while those with $N\leq 3$ exhibit the strain-softening typical of convex particles. While the stress $\sigma$ tends to get larger with increasing $N$, the curves in Fig. 3 are not monotonically increasing with $N$, indicating a large run-to-run variation ($\sim 30\%$) compared to any trends in $N$ for $N\geq 9$.
@@ III.2 Entangled area
To develop a quantitative model to predict the probability of entanglements, we first measure the free enclosed area $A$ enclosed by each chain which is available for other chains to pass through. This is calculated from the x-ray tomography data as the area in the entanglement manifold that is at least 1 bead diameter away from a bead in the entangling chain, as illustrated in Fig. 4b. This condition accounts for the fact that the finite bead size excludes the center of other beads from overlapping with the beads in the entangling chain, a.k.a. an excluded volume effect. The average values for each contour length $m$ are shown in Fig. 5, shown in units of bead diameter squared. The shorter contour lengths have very little free enclosed area $A$ because the chains are short compared to the persistence length or minimum loop size and tend to be straighter. For contour lengths of $m\leq 3$, the finite bead size completely excludes chains from passing through the enclosed area for this minimum loop size. This explains why no entanglements were observed for $m\leq 3$ Brown et al. (2012). Since chains that are longer can be bent into larger arcs, the free enclosed area tends to increase with $m$, similar to the increase in number of entanglements with $m$ Brown et al. (2012).
Figure 6: Number of entanglements $n$ per enclosure as a function of free enclosed area $A$. The points correspond to different contour lengths $m$, using the same data as Fig. 5. Circles: chain length $N=20$. Diamonds: $N=10$. Open symbols: $\gamma=0$. Filled symbols: $\gamma=0.2$. Crossed symbols: plotted against free enclosed area $A$. Non-crossed symbols: plotted against effective free enclosed area $A_{eff}$ including constraints in the third dimension. Lines: linear fits to $N=10$ data. The near proportionality between $n$ and $A$ or $A_{eff}$ suggests that the probability of entanglement is determined by random arrangements with excluded volume effects, so the average number of entanglements and whether or not there are likely to be system-filling clusters of entanglements can be predicted from particle shapes in the packing.
@@
The number of entanglements $n$ is predicted based on random arrangements with excluded volume effects using a model calculation that was originally developed for staple packings Gravish et al. (2012). It involves equating two calculations of volume of entangled particles inside the volume made by an entanglement manifold of thickness $\delta$ – in this case we set $\delta$ equal to one bead diameter $a$. The average volume of chains in an entangled volume $V_{ent}$ can be estimated as the average number of entanglements $n$ times the volume per unit length of chain taken up by an entangled bead $\pi a^{3}/6\bar{d}$ (with average center-to-center spacing $\bar{d}$) times the length $\delta/\cos\Theta$ in the entangled plane of thickness $\delta$, where $\Theta$ is the angle of the entangled chain relative to a line perpendicular to the manifold (see illustration in Fig. 22 of Appendix E). This results in $V_{ent}=n\pi a^{3}\delta/6\bar{d}\cos\Theta$. This is balanced with a prediction of $V_{ent}$ based on the assumption that particle positions are random so the volume of entangled particles is equal to the packing fraction times the free volume available inside the entanglement. This is calculated as $V_{ent}=\phi A(\Theta,\Phi)\delta$, where $A(\Theta,\Phi)$ is a modification of $A$ which also accounts for the three-dimensional excluded volume from different possible angles $\Theta$ and $\Phi$ around this axis. Equating the two calculations of $V_{ent}$ results in $n=6\bar{d}\phi\langle A(\Theta,\Phi)\cos\Theta\rangle/\pi a^{3}=0.93\langle A(\Theta,\Phi)\cos\Theta\rangle/a^{2}$ for $\phi=0.44\pm 0.03$ (see Appendix D for measurements of $\phi$) and average center-to-center distance $\bar{d}/a=1.11$ (measured from x-ray tomography). The effective free enclosed area $A_{eff}=\langle A(\Theta,\Phi)\cos\Theta\rangle$ accounts for excluded volume effects in three dimensions, as well as the larger cross sections of entangled chains at various angles $\Theta$. $A_{eff}/A$ corresponds to the parameter $1/\alpha$ introduced as a fit parameter by Gravish et al. (2012). The calculation of $A_{eff}$ from our x-ray tomography data is explained in Appendix E. The number of entanglements $n$ as a function of $A_{eff}$ calculated from x-ray tomography data is shown in Fig. 6. This also appears nearly linear, as there is nearly a constant ratio $A_{eff}/A=0.27\pm 0.01$ in the strong strain-stiffening range with $m\geq 8$ (see Appendix E). For the $N=10$ data in Fig. 6 with percentage errors, a linear fit yields $n=(0.87\pm 0.14)A_{eff}/a^{2}$ and a coefficient of determination of 0.92, confirming a nearly linear trend as predicted, with a coefficient consistent with the predicted value of $n=0.93A_{eff}/a^{2}$. This confirms that the assumption of random positions and orientations with excluded volume effects can relate the number of entanglements $n$ to the free enclosed area $A_{eff}$ if the shape of the particles in the packing is known, at least for $N=10$. The measured ratio $n/A_{eff}$ decreases for $N=20$ and for increasing $m$, suggesting that longer chains have additional constraints that reduce the number of entanglements, for example due to the maximum bend angle of chains which puts constraints on particle positions in the same chain. Since $n/A$ is more linear than $n/A_{eff}$ for small $m$, the larger $n/A_{eff}$ for small $m$ with the assumption of random particle orientations $\Theta$ in $A_{eff}$ indicates the angles $\Theta$ are more perpendicular to the entangled plane than random orientations, which could be a result of contact forces tending to push entangled chains into the small range of angles that they will fit.